OpenStudy (anonymous):

which of the following represents a range of the function f(x)=-4x+1 for the domain -5,-1,0,2,10

OpenStudy (anonymous):

a. (-39,-7,1,5,21) b.(-41,-9,1,5,21) c.(-19,-3,1,5,21) d.(-41,-9,-3,-19)

OpenStudy (texaschic101):

range is all the y values and domain is all the x values. f(x) = -4x + 1 f(-5) = -4(-5) + 1 f(-5) = 20 + 1 f(-5) = 21 f(-1) = -4(-1) + 1 f(-1) = 4 + 1 = 5 f(0) = -4(0) + 1 f(0) = 1 f(2) = -4(2) + 1 f(2) = -8 + 1 f(2) = - 7 f(10) = -4(10) + 1 f(10) = -40 + 1 f(10) = -39 so we have : -39,-7,1,5,21

OpenStudy (anonymous):

Thanks you so much! :)

OpenStudy (texaschic101):

YW :)